OpenStudy (anonymous):

Which of the following is a counter example of the statement? If 2n + 4 < 20, then n < 6

OpenStudy (anonymous):

7 would work

OpenStudy (anonymous):

n<8

OpenStudy (anonymous):

If n = 5, 2n + 4 < 20 is true. If n = 7, 2n + 4 < 20 is true. If n = 10, 2n + 4 < 20 is false. If n= 20, 2n + 4 < 20 is false. my options